OpenStudy (anonymous):

Fanning + Medaling: ( 1/ sqrt1 +t) -1 Lim ______________ t->o t

OpenStudy (anonymous):

\[\lim_{t \rightarrow 0}\frac{ \frac{ 1 }{ \sqrt{1} + t } -1 }{ t }\]